# Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator

### v1

@@ II Discussion
Our results provide a simple and elegant solution to a problem that has eluded researchers for nearly half a century. Through exact results and detailed numerical simulations, we demonstrate that the long-sought-after RVB spin liquid state exhibiting spin-charge separation emerges naturally as ground state of a large-$U$ Hubbard Hamiltonian on lattices of corner-sharing tetrahedra. The key aspect of this emergent behaviour is the kinetic energy frustration of carriers that result from doping the Mott state.
@@
The exact results found in this work about the ground state wave function of the system, both in the thermodynamic limit as well as for finite size, raise interesting questions about the nature of the excitations, and in particular the statistics of the holons, which shall be left to future work.
@@
The eminently neat and realistic nature of the Hamiltonian opens the door to further theoretical investigations into the physics of RVB phases, possibly also with the help of implementations on synthetic quantum platforms [Trabesinger2012, Georgescu2014] (note in particular how the leading RVB liquid behaviour is already apparent in relatively small systems of 16 sites). The rare availability of a seemingly exact single-hole ground state wave function for systems of any size provides uncommon ability to compare theory and simulator experiments.
[…]
@@
Our results apply to corner-sharing tetrahedral lattices in general, including the two-dimensional checkerboard lattice. While less common than the pyrochlore lattice in real materials, one wonders whether such geometry could emerge in quasi-2D square lattice compounds, possibly mediated by inter-layer atoms; or perhaps in 2D materials and synthetic quantum platforms. From a theoretical perspective, it would be interesting to investigate in greater detail the robustness of the RVB liquid behaviour and spin-charge separation to an asymmetry between the nearest-neighbour and next-nearest-neighbour (diagonal) hopping terms on the checkerboard lattice.
@@
As speculated earlier, the RVB singlet state in our system could exhibit an emergent $U(1)$ gauge symmetry with power-law singlet correlators and gapless (photon) excitations (in 3D) [Baskaran1988, Read1989, Fradkin1990]. In presence of itinerant (fermionic [Read1989b]) holons, this is reminiscent of models used to investigate Fermi Liquid Star (FL*) behaviour [Senthil2003], and may provide an additional experimental avenue to explore it.
@@
In closing, our work proposes kinetic energy frustration of doped Mott insulators as a novel route to realising robust topologically ordered states, in contrast to the conventional approach based on geometrically frustrated interactions. This holds great promise for the theoretical and experimental discovery of new emergent states of matter. As for a potential connection between kinetically frustrated RVB liquids and superconductivity [Anderson1987], this shall remain an open question for future work.
III Online Methods
We studied a 16-site system consisting of a single conventional cubic unit cell of the pyrochlore lattice (see Supp. Mat. Note C for illustrations) using ED with Lanczos to compute the lowest eigenvalues and corresponding eigenstates in each magnetisation sector. We also considered a 32-site system consisting of a $2\times 2\times 2$ cell constructed from the fcc primitive lattice vectors. For this system size, we used two-site DMRG with an energy convergence criterion of $10^{-10}$ and a maximum bond dimension of $12800$. We imposed periodic boundary conditions, creating a 1D snake path of the system following the numbering shown in the Supp. Mat. Note C. The choice of snake path seems to have little impact on the convergence for our highly entangled 3D system. The DMRG calculations were performed using the TeNPy Library (version 0.10.0) [tenpy].
A single DMRG calculation gives one ground state. When the ground state is degenerate, we biased each ground state and compute correlations as an average over all ground states. In the case of pure hopping ($J=0$), the full ground state manifold was found by pinning the unpaired spin by adding an on-site magnetic field. This gives an overcomplete set of ground states, and a linearly independent set of ground states is then found via Gram-Schmidt decomposition.
A similar approach was deployed to study our system on the checkerboard lattice, using ED for 16 sites (1 and 2 holes) and 24 sites (1 hole), and DMRG for 24 sites (2 holes).
@@ Acknowledgements
The project was conceived by C.C. and C.G., and all authors contributed to its development, with specific focus by C.G. on the numerical efforts; by C.D.B. on the analytical proofs; and by P.A.G. and D.P. on the experimental component. All authors contributed to writing the manuscript.
@@
Supplementary Material for:
“Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator”
Cecilie Glittum,1 Antonio Štrkalj,1,2 Dharmalingam Prabhakaran,3 Paul A. Goddard,4 Cristian D. Batista,5,6 and Claudio Castelnovo1
1T.C.M. Group, Cavendish Laboratory, JJ Thomson Avenue, Cambridge CB3 0HE, United Kingdom
2Department of Physics, Faculty of Science, University of Zagreb, Bijenička c. 32, 10000 Zagreb, Croatia
3Department of Physics, University of Oxford, Clarendon Laboratory, Parks Road, Oxford OX1 3PU, UK
4Department of Physics, University of Warwick, Coventry CV4 7AL, United Kingdom
5Department of Physics and Astronomy, University of Tennessee, Knoxville, TN 37996, USA
6Neutron Scattering Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831, USA
@@ A.1 Ground state energy
Since the Hamiltonian $\hat{\mathcal{H}}$ includes only nearest neighbour hopping terms, it can be decomposed as a sum of Hamiltonians ^$\hat{\mathcal{H}}^{\alpha}$ acting on each tetrahedron $\alpha$:
^$\hat{\mathcal{H}}=\sum_{\alpha}\hat{\mathcal{H}}^{\alpha}\,,$ (3)
where the dual lattice index $\alpha$ labels the different tetrahedra and ^$\hat{\mathcal{H}}^{\alpha}$ includes only the six bonds of the tetrahedron $\alpha$.
@@
Eq. (9) implies that the ground state energy of a single tetrahedron is lower than $-2t$. However, the ground state energy of a single tetrahedron is exactly $-2t$, which implies that $E_{g}$ cannot be lower than $-4t$.
For completeness, let us close by showing explicitly that the single tetrahedron GS energy is $-2t$. If we consider a system consisting of four fully connected sites, labelled $1,2,3,4$, with one holon and three spins, we can define the wave functions
@@
with $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. Note that $\hat{d}^{\dagger}_{ij}=\hat{d}^{\dagger}_{ji}$. The state $|\Psi^{\rm tet}_{i\sigma}\rangle$ has a static spin $\sigma$ on site $i$, while the holon and a singlet resonate on the remaining three sites. Consequently, the total spin of the state is $\sigma$. Applying the single tetrahedron Hamiltonian to this state, we find that hopping to site $i$ indeed is cancelled, and the states are eigenstates with energy $-2t$. Note that the state where site $4$ has a static spin is a superposition of the three states in Eq. (10): $|\Psi^{\rm tet}_{4\sigma}\rangle=\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{4\sigma}\left(\hat{d}^{\dagger}_{12}+\hat{d}^{\dagger}_{23}+\hat{d}^{\dagger}_{31}\right)|0\rangle=-(|\Psi^{\rm tet}_{1\sigma}\rangle+|\Psi^{\rm tet}_{2\sigma}\rangle+|\Psi^{\rm tet}_{3\sigma}\rangle)$.
@@ A.2 Ground state wave function
^$\hat{\mathcal{H}}^{\alpha}|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle=-2t\,|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle\,,$ (23)
@@
Since we have shown that $-4t$ is a strict lower bound for the energy, Eq. (24) implies that $|\Psi_{\rm RVB}\rangle$ is an exact ground state of $\hat{\mathcal{H}}$. Eq. (23) clarifies the name ‘passive’ for the site ${\bm{r}}$: since $|\tilde{\Psi}^{\alpha}_{\bm{r}}\rangle$ is an eigenstate of ^$\hat{\mathcal{H}}^{\alpha}$, the holon never visits that site because of the destructive interference produced by the linear superposition (20).
@@ A.4 Proposed finite-size ground state wave function
From numerics (discussed in Supp. Mat. Notes C and D), we find that the one-hole GS energy $-4t$ can be met in finite size systems and even when the unpaired spin is fixed at a given site; in the latter case the GS is unique. This allows us to conjecture a form for the wave function that is given by the equal amplitude superposition of all dimer-singlet states that have a spin fixed and a delocalised holon, subject to the constraint of no more than one dimer per tetrahedron. Remarkably, we find that the conjectured state is identical within numerical accuracy to the ground state found for the 16-site system using ED, and to the 32-site system using DMRG. We also find this to be the case for the two-holes ground state (i.e., the equal amplitude superposition of all dimer-singlet states that have two delocalised holons, subject to the constraint of no more than one dimer per tetrahedron). We therefore propose that such GS wave functions are in fact exact for all system sizes.
@@ A.5 Vanishing wave function in the stoichiometric limit
More recently, Haerter and Shastry [Haerter2005] pointed out that the same kinetic mechanism can give rise to strikingly different spin correlations – namely, weak metallic antiferromagnetism – if the electronic lattice is frustrated; this occurs, for instance, when the electron hopping amplitude is negative, and the elementary loops on the lattice encompass an odd number of bonds. This counterpart to the Nagaoka effect was further studied in Refs. Sposetti2014, Lisandrini2017 and later dubbed the ‘counter-Nagaoka effect’ [Zhang2018].
The nature of the state on the frustrated triangular lattice was investigated numerically [Sposetti2014, Lisandrini2017, Morera2021] and shown to exhibit correlations that depart significantly from the behaviour expected for the quantum Heisenberg antiferromagnet, towards a more classical nature (although not entirely: a small deviation from classical correlations is observed [Sposetti2014]). This work also demonstrated that the kinetic energy mechanism was akin to a finite spin interaction term proportional to the hole density (namely, $J_{t}\sim t/N$, where $N$ is the number of lattice sites, in the presence of a single hole).
The effect has been recently investigated in ladders in Ref. Morera2021; and on a Husimi cactus of triangles (akin locally to a kagome lattice) where it was shown to form a regular valence bond pattern [Kim2023].
@@
Figure S1: 16-site pyrochlore system, constructed from a single cubic unit cell with lattice vectors ${\bm{a}}_{1}=(1,0,0)$, ${\bm{a}}_{2}=(0,1,0)$ and ${\bm{a}}_{3}=(0,0,1)$.
Figure S2: 32-site pyrochlore system, constructed from a $2\times 2\times 2$ lattice of the ffc unit cell with lattice vectors ${\bm{a}}_{1}=(0,1/2,1/2)$, ${\bm{a}}_{2}=(1/2,0,1/2)$ and ${\bm{a}}_{3}=(1/2,1/2,0)$. Each unit cell has four sites (a single tetrahedron).
@@
^$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{P}_{i\sigma}\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}\hat{P}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{{\bm{S}}}_{i}\cdot\hat{{\bm{S}}}_{j}\,,$ (25)
where once again $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the fermionic creation (annihilation) operator at site $i$; and $\hat{{\bm{S}}}_{i}$ is the corresponding spin operator, $\hat{{\bm{S}}}_{i}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$.
@@
We use two-site DMRG to study the 32-site system, with an energy convergence criterion of $10^{-10}$ and a maximum bond dimension up to $12800$. Using periodic boundary conditions, we create a 1D snake path of the system that follows the numbering in Fig. S2. The choice of snake path seems to have little impact on the convergence for our highly entangled 3D system. The DMRG calculations were performed using the TeNPy Library (version 0.10.0) [tenpy].
@@
As mentioned in the main text, our results hold for lattices of corner sharing tetrahedra and therefore also for the 2D checkerboard lattice. Indeed, the 16-site unit cell of the 2D checkerboard lattice is identical to the one for the pyrochlore lattice. In our work, we further consider a 24-site checkerboard lattice using ED (1 hole) and DMRG (2 holes).
@@
In the first instance, we consider the case of non-interacting spins ($t=1$, $J=0$) at zero temperature in the presence of a single hole. For all three system sizes, we find that the ground state energy bound is met at $-4t$. For 16 and 32 sites, the GS degeneracy is $18$ and $34$, respectively.
@@
The holon density shows weak repulsive dependence on the distance from the fixed spin, as illustrated in Fig. 2 in the main text. Correspondingly, the expectation value of the hole hopping operator has a suitably inverse dependence, leading to the exact Hamiltonian eigenvalue $-4t$. This is confirmed by computing the expectation value $\langle\sum_{\sigma,j}\hat{c}^{\dagger}_{i,\sigma}\hat{c}_{j,\sigma}\rangle$, where $j$ runs over all nearest-neighbour sites of $i$. Within numerical accuracy, we find indeed that this expectation value is precisely $4\langle 1-\hat{n}_{i}\rangle$ for all sites $i$, where $\langle 1-\hat{n}_{i}\rangle$ is the holon density at site $i$.
@@
We find within numerical accuracy that the GS wave function (with a fixed spin) from 16-site ED and 32-site DMRG for the pyrochlore lattice and 16- and 24-site ED for the checkerboard lattice is identical to a state constructed as the equal amplitude superposition of dimer-singlet states with a holon and the further constraint of no more than one dimer per tetrahedron.
@@
Figure S3: Spin structure factors of the 32-site systems. Top: Doping-induced, $1$ hole ($t=1$ and $J=0$). Middle: Quantum Heisenberg model (QHM). Bottom: Classical Heisenberg model (CHM). Both Heisenberg models have no holes and their correlations are correspondingly scaled by $(N-1)/N$. Accessible momenta are marked by grey pluses.
@@
We compute the GS in the presence of two holes and study the GS spin and holon correlations for $t=1$ and $J=0$. We find that the GS is unique with energy $-8t$ in the $S^{z}=0$ sector for both the 16-site and 32-site systems, which is exactly twice the GS energy of a single hole.
@@
Figure S4: Dependence of the ground state spin structure factor (for all individual reciprocal lattice points) on $J/t$ ($t>0$) for the 16-site system with two holes. Dashed horizontal lines show the $t=0$ limits, for reference. The shaded region denotes correlations that are antiferromagnetic in spite of a ferromagnetic interaction $J$. The dashed vertical line shows $(J/t)_{c}\simeq-0.173$ for holon density $1/16$.
Next, we study the effect of having a finite $J$ when two holes are present in the 16-site system. In Fig. S4 we show the GS spin structure factor as a function of $J/t$ with $t=1$. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.173$ (hole density $12.5$%). This shows that also in the presence of two holes, quantum spin liquid behaviour can be induced over magnetic ordering by frustrated hole doping.

### v2

@@ II Discussion
Our results provide a simple and elegant solution to a problem that has eluded researchers for nearly half a century. Through exact results and detailed numerical simulations, we demonstrate that the long-sought-after RVB spin liquid state exhibiting spin-charge separation emerges naturally as ground state of a large-$U$ Hubbard Hamiltonian on lattices of corner-sharing tetrahedra, upon dilute-doping away from the Mott insulating state. The key aspect of this emergent behaviour is the kinetic energy frustration of carriers that result from doping the Mott state.
@@
The eminently neat and realistic nature of the Hamiltonian opens the door to further theoretical investigations into the physics of RVB phases, possibly also with the help of implementations on synthetic quantum platforms [Trabesinger2012, Georgescu2014] (note in particular how the leading RVB liquid behaviour is already apparent in relatively small systems of 16 sites). The rare availability of an exact single-hole and two-hole ground state wave function for systems of any size provides uncommon ability to compare theory and simulator experiments.
It should also be possible to test our results in real materials. In the first instance, an experimental comparison between the effects of hole and electron doping, i.e., switching between counter-Nagaoka antiferromagnetic correlations and Nagaoka ferromagnetic ones [Ciociaro2023] (as shown in Fig. 3), could be used to check if a given Mott insulator is indeed in the large-$U$ regime relevant to this work. The dynamical spin structure factor $S({\bm{Q}},\omega)$ of the predicted RVB state can be experimentally accessed via neutron scattering, as a function of reciprocal lattice vector $\bm{Q}$ and angular frequency $\omega$. The low-energy region, in particular, ought to include a spinon continuum, flat in the infinite-$U$ limit and broadened by $J$, which could be used to measure spin interactions. The spin-spin correlation length can be directly extracted from the static spin structure factor $S({\bm{Q}})=\int d\omega S({\bm{Q}},\omega)$. Spin-charge separation can be indirectly measured via photo-emission experiments, which will reveal a vanishing weight of the quasi-particle peak and can also be used to extract the hopping amplitude $t$. Finally, in presence of interactions (e.g., exchange $J$) the undoped Mott insulating phase is likely to exhibit magnetic properties that are notably different from those of the RVB liquid phase. A transition between the two phases as a function of doping could be detected via signatures in thermodynamic properties, such as specific heat and susceptibility, as well as in neutron scattering or muon-spin rotation measurements, raising the tantalising prospect of observing hole-induced quantum spin liquid behaviour. If the interactions are ferromagnetic, one could envisage looking for polaronic behaviour and possibly magnonic Cooper pairs (bound states of two fermions and one magnon) in the ferromagnetic phase close to the transition to the RVB state [Zhang2018, nazaryan2024].
Candidate materials that realise large-$U$-Hubbard-model Mott insulators on lattices of corner-sharing tetrahedra are not commonplace. One can nonetheless identify families of compounds which, while they do not immediately map onto our model, provide promising avenues for further investigation and the possible realisation of the RVB liquid behaviour presented in our work. One such family of compounds, the $5d$ transition metal pyrochlore oxides, has been extensively studied and is known to form regular pyrochlore lattices. Rare-earth iridates and ruthenates, RE2TM2O7, host $5d^{5}$ TM4+ ions with an effective angular momentum $J_{\rm eff}=1/2$ ground-state doublet. Holons can be introduced by doping on the non-magnetic RE site, thus effectively replacing TM4+ with TM5+, $J_{\rm eff}=0$ ions. Some doping experiments of these materials have already been performed, including measurements on polycrystalline Y2-xCaxIr2O7 for x=0x=00.20.2 [Fukazawa2002, Zhu2014] where a marked effect on the magnetic and transport properties was observed. However, the polycrystalline nature of the samples used in these studies seems to have led to confusing and contradictory results (see Supp. Mat. Note G for additional details). Recent progress in producing high-quality single crystals of this material family (see, e.g., Refs. Ishikawa2012, Cathelin2020, Uehara2022 and inset to Fig. 1), could provide the opportunity to understand these changes and perhaps observe evidence for the predictions discussed above. (See Supp. Mat. Note G for a more detailed review of the materials that could potentially be investigated, the methods by which they could be grown, and possible experimental routes to measure a doping-induced order-to-QSL transition.) Having said this, certain caveats should be flagged regarding the applicability of our theoretical results to these $5d$ compounds: they tend to be weak insulators, and they display strong spin-orbit effects; hence they are less likely to be well-described by models of isotropic hopping. Further theoretical work is needed to study the tolerance of the original model to these departures. Another family of compounds involves $3d$ transition metals, which offer large $U/t$ ratios and (owing to weak spin-orbit coupling) near-isotropic hopping intrinsic to our model. Investigating pyrochlore lattices of spin-1/2 Cu(II) is particularly promising for two reasons: they are strong Mott insulators with a large $U/t$ ratio, and their spin-orbit coupling is substantially weaker than that of 4dd and 5dd elements. However, only a few suitable systems have been identified to date (see e.g., Refs. Pinsard2004, Wills2004, Uematsu2007, Kawabata2007, Nishiyama2011) and these often exhibit distortions from the ideal pyrochlore lattice. The impact of such distortions on our model would need to be carefully considered.
@@
Our results apply to corner-sharing tetrahedral lattices in general, including the two-dimensional checkerboard lattice [Poilblanc2004] (see Supp. Mat. Note F for a study of 16- to 48-site checkerboard systems). While less common than the pyrochlore lattice in real materials, one wonders whether such geometry could emerge in quasi-2D square lattice compounds, possibly mediated by inter-layer atoms; or perhaps in 2D materials and synthetic quantum platforms. From a theoretical perspective, it would be interesting to investigate in greater detail the robustness of the RVB liquid behaviour and spin-charge separation to an asymmetry between the nearest-neighbour and next-nearest-neighbour (diagonal) hopping terms on the checkerboard lattice. It would also be interesting to study the effects of non-magnetic vacancy disorder, which has been recently argued to destabilise some but not all short-range RVB liquid phases in 2D, depending on the lattice [Ansari2024].
@@
As speculated earlier, the RVB singlet state in our system could exhibit an emergent $U(1)$ gauge symmetry with power-law singlet correlators and gapless (photon) excitations (in three dimensions) [Baskaran1988, Read1989, Fradkin1990]. In presence of itinerant (fermionic [Read1989b]) holons, this is reminiscent of models used to investigate Fermi Liquid Star (FL*) behaviour [Senthil2003], and may provide an additional experimental avenue to explore it.
@@
In closing, our work proposes kinetic energy frustration of doped Mott insulators as a route to realising robust topologically ordered states, in contrast to the conventional approach based on geometrically frustrated interactions. This holds great promise for the theoretical and experimental discovery of emergent states of matter. As for a potential connection between kinetically frustrated RVB liquids and superconductivity [Anderson1987], this shall remain an open question for future work.
@@ Acknowledgements
The project was conceived by C.C. and C.G., and all authors contributed to its development, with specific focus by C.G. on the numerical efforts with A.Š. as discussion partner, by C.D.B. on the analytical proofs, and by P.A.G. and D.P. on the experimental component. All authors contributed to writing the manuscript.
@@
Supplementary Material Table of Content
Supp. Mat. Note A: Proofs
Supp. Mat. Note B: The counter-Nagaoka effect
Supp. Mat. Note C: Models and methods
Supp. Mat. Note D: Doping and correlations
Supp. Mat. Note E: Two holes
Supp. Mat. Note F: Checkerboard
Supp. Mat. Note G: Experiments
@@ A.1 Ground state energy
Since the Hamiltonian $\hat{\mathcal{H}}$ includes only nearest neighbour hopping terms, it can be decomposed as a sum of Hamiltonians $\hat{\mathcal{H}}^{\alpha}$ acting on each tetrahedron $\alpha$:
$\hat{\mathcal{H}}=\sum_{\alpha}\hat{\mathcal{H}}^{\alpha}\,$, (3)
where the dual lattice index $\alpha$ labels the different tetrahedra and $\hat{\mathcal{H}}^{\alpha}$ includes only the six bonds of the tetrahedron $\alpha$.
@@
Eq. (9) implies that the ground state energy of a single tetrahedron is lower than 2t-2t. However, the ground state energy of a single tetrahedron is exactly 2t-2t, which implies that $E_{g}$ cannot be lower than 4t-4t.
For completeness, let us close by showing explicitly that the single tetrahedron GS energy is 2t-2t. If we consider a system consisting of four fully connected sites, labelled $1,2,3,4$, with one holon and three spins, we can define the wave functions
@@
with $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. Note that $\hat{d}^{\dagger}_{ij}=\hat{d}^{\dagger}_{ji}$. The state $|\Psi^{\rm tet}_{i\sigma}\rangle$ has a static spin $\sigma$ on site $i$, while the holon and a singlet resonate on the remaining three sites. Consequently, the total spin of the state is $\sigma$. Applying the single tetrahedron Hamiltonian to this state, we find that hopping to site $i$ indeed is cancelled, and the states are eigenstates with energy 2t-2t. Note that the state where site $4$ has a static spin is a superposition of the three states in Eq. (10): $|\Psi^{\rm tet}_{4\sigma}\rangle=\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{4\sigma}\left(\hat{d}^{\dagger}_{12}+\hat{d}^{\dagger}_{23}+\hat{d}^{\dagger}_{31}\right)|0\rangle=-(|\Psi^{\rm tet}_{1\sigma}\rangle+|\Psi^{\rm tet}_{2\sigma}\rangle+|\Psi^{\rm tet}_{3\sigma}\rangle)$.
@@ A.2 Ground state wave function
$\hat{\mathcal{H}}^{\alpha}|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle=-2t\,|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle\,$, (23)
@@
Since we have shown that 4t-4t is a strict lower bound for the energy, Eq. (24) implies that $|\Psi_{\rm RVB}\rangle$ is an exact ground state of $\hat{\mathcal{H}}$. Eq. (23) clarifies the name ‘passive’ for the site $\bm{r}$}: since $|\tilde{\Psi}^{\alpha}_{\bm{r}}\rangle$ is an eigenstate of $\hat{\mathcal{H}}^{\alpha}$, the holon never visits that site because of the destructive interference produced by the linear superposition (20).
@@ A.4 Proposed finite-size ground state wave function
From numerics (discussed in Supp. Mat. Notes C and D), we find that the one-hole GS energy 4t-4t can be met in finite size systems and even when the unpaired spin is fixed at a given site; in the latter case the GS is unique. This allows us to conjecture a form for the wave function that is given by the equal amplitude superposition of all dimer-singlet states that have a spin fixed and a delocalised holon, subject to the constraint of no more than one dimer per tetrahedron. Remarkably, we find that the conjectured state is identical within numerical accuracy to the ground state found for the 16-site system using ED, and to the 32-site system using DMRG. We also find this to be the case for the two-holes ground state (i.e., the equal amplitude superposition of all dimer-singlet states that have two delocalised holons, subject to the constraint of no more than one dimer per tetrahedron). We therefore propose that such GS wave functions are in fact exact for all system sizes.
@@ A.5 Vanishing wave function in the stoichiometric limit
More recently, Haerter and Shastry [Haerter2005] pointed out that the same kinetic mechanism can give rise to strikingly different spin correlations – namely, weak metallic antiferromagnetism – if the electronic lattice is frustrated; this occurs, for instance, when the electron hopping amplitude is negative, and the elementary loops on the lattice encompass an odd number of bonds (see also Ref. Poilblanc2004 for earlier results hinting at this phenomenon). This counterpart to the Nagaoka effect was further studied in Refs. Sposetti2014, Lisandrini2017 and later dubbed the ‘counter-Nagaoka effect’ [Zhang2018].
The nature of the state on the frustrated triangular lattice was investigated numerically [Sposetti2014, Lisandrini2017, Morera2024] and shown to exhibit correlations that depart significantly from the behaviour expected for the quantum Heisenberg antiferromagnet, towards a more classical nature (although not entirely: a small deviation from classical correlations is observed [Sposetti2014]). This work also demonstrated that the kinetic energy mechanism was akin to a finite spin interaction term proportional to the hole density (namely, $J_{t}\sim t/N$, where $N$ is the number of lattice sites, in the presence of a single hole).
The effect has been recently investigated in ladders in Ref. Morera2024; and on a Husimi cactus of triangles (akin locally to a kagome lattice) where it was shown to form a regular valence bond pattern [Kim2023].
@@
Figure S1: 16-site pyrochlore system. This is constructed from a single cubic unit cell with lattice vectors ${\bm{a}}_{1}=(1,0,0)$, ${\bm{a}}_{2}=(0,1,0)$ and ${\bm{a}}_{3}=(0,0,1)$.
Figure S2: 32-site pyrochlore system. This is constructed from a $2\times 2\times 2$ lattice of the ffc unit cell with lattice vectors ${\bm{a}}_{1}=(0,1/2,1/2)$, ${\bm{a}}_{2}=(1/2,0,1/2)$ and ${\bm{a}}_{3}=(1/2,1/2,0)$. Each unit cell has four sites (a single tetrahedron).
@@
$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{P}_{i\sigma}\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}\hat{P}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{{\bm{S}}}_{i}\cdot\hat{{\bm{S}}}_{j}\,$, (25)
where once again $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the fermionic creation (annihilation) operator at site $i$; and $\hat{{\bm{S}}}_{i}$ is the corresponding spin operator, $\hat{{\bm{S}}}_{i}=\tfrac{1}{2}\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$.
@@
Figure S3: Energy as a function of the inverse bond dimension for the 32-site pyrochlore system. The bond dimensions used are $\chi=\{800,1000,1600,3200,6400,12800\}$.
We use two-site DMRG to study the 32-site system, with an energy convergence criterion of $10^{-10}$ and a maximum bond dimension up to $12800$. This is sufficient to reach convergence, as shown in Fig. S3. Using periodic boundary conditions, we create a one-dimensional snake path of the system that follows the numbering in Fig. S2. The choice of snake path seems to have little impact on the convergence for our highly entangled 3D system. The DMRG calculations were performed using the TeNPy Library (version 0.10.0) [tenpy].
@@
As mentioned in the main text, our results hold for lattices of corner sharing tetrahedra and therefore also for the 2D checkerboard lattice. Indeed, the 16-site unit cell of the 2D checkerboard lattice ($4\times 2\times 2$) is identical to the one for the pyrochlore lattice ($16\times 1\times 1\times 1$). In our work, we further consider a 24-site ($4\times 3\times 2$) checkerboard lattice using ED (1 hole) and DMRG (2 holes) as well as 36- ($4\times 3\times 3$) and 48-site ($4\times 4\times 3$) checkerboard lattices with 1 and 2 holes using DMRG.
@@
In the first instance, we consider the case of non-interacting spins ($t=1$, $J=0$) at zero temperature in the presence of a single hole. For all system sizes, we find that the ground state energy bound is met at 4t-4t. For 16 and 32 sites, the GS degeneracy is $18$ and $34$, respectively.
@@
The holon density shows weak repulsive dependence on the distance from the fixed spin, as illustrated in Fig. 2 in the main text. Correspondingly, the expectation value of the hole hopping operator has a suitably inverse dependence, leading to the exact Hamiltonian eigenvalue 4t-4t. This is confirmed by computing the expectation value $\langle\sum_{\sigma,j}\hat{c}^{\dagger}_{i,\sigma}\hat{c}_{j,\sigma}\rangle$, where $j$ runs over all nearest-neighbour sites of $i$. Within numerical accuracy, we find indeed that this expectation value is precisely $4\langle 1-\hat{n}_{i}\rangle$ for all sites $i$, where $\langle 1-\hat{n}_{i}\rangle$ is the holon density at site $i$.
@@
We find within numerical accuracy that the GS wave function (with a fixed spin) from 16-site ED and 32-site DMRG for the pyrochlore lattice, and 16- and 24-site ED for the checkerboard lattice , is identical to a state constructed as the equal amplitude superposition of dimer-singlet states with a holon and the further constraint of no more than one dimer per tetrahedron.
@@
Figure S4: Spin structure factors of the 32-site systems. Top, doping-induced, 11 hole ($t=1$ and $J=0$). Middle, quantum Heisenberg model (QHM). Bottom, classical Heisenberg model (CHM). Both Heisenberg models have no holes and their correlations are correspondingly scaled by (N1)/N(N-1)/N. Accessible momenta are marked by grey pluses.
@@
We compute the GS in the presence of two holes and study the GS spin and holon correlations for $t=1$ and $J=0$. We find that the GS is unique with energy 8t-8t in the $S^{z}=0$ sector for both the 16-site and 32-site systems, which is exactly twice the GS energy of a single hole.
@@
Figure S5: Dependence of the ground state spin structure factor on $J/t$ ($t>0$). This is shown for all individual reciprocal lattice points for the $16$-site system with two holes (density $1/8$). Dashed horizontal lines show the $t=0$ limits, for reference. The shaded region denotes correlations that are antiferromagnetic in spite of a ferromagnetic interaction $J$, extending down to $(J/t)_{c}\simeq-0.173$. For comparison, the dashed vertical line inside the shaded region shows the corresponding extent for hole density $1/16$, $(J/t)_{c}\simeq-0.052$ (see Fig. 3 in the main text).
Next, we study the effect of having a finite $J$ when two holes are present in the 16-site system. In Fig. S5 we show the GS spin structure factor as a function of $J/t$ with $t=1$. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.173$ (hole density 12.512.5%). This shows that also in the presence of two holes, quantum spin liquid behaviour can be induced over magnetic ordering by frustrated hole doping.